OpenStudy (anonymous):

How do I find the quotient of 2 2/3 and 1/4?

OpenStudy (anonymous):

fist you must know that quotient means result of a division problem so we need to divide 2 2/3 by 1/4. Do this by converting to improper fractions \[\frac{ 8 }{ 3 }\div \frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

dividing by a fraction is the same thing as multiplying by its reciprocal\[\frac{ 8 }{ 3 }\div \frac{ 1 }{ 4 }=\frac{ 8 }{ 3 }\times \frac{ 4 }{ 1 }\]

OpenStudy (anonymous):

So the answer would be 2/3

OpenStudy (anonymous):

\[\frac{ 8 }{ 3 }\times \frac{ 4 }{ 1 } = \frac{ 8\times 4 }{ 3 \times 1 } = \frac{ ? }{ ? } \]